OpenStudy (jkbo):

Solve the system. -6 = 3x - 6y 4x = 4 + 5y A.no solution B.(6, 4) C.infinite solutions D.(6, -4)

OpenStudy (anonymous):

B.

OpenStudy (jkbo):

@UnixEditz can you shoq mw hoq you got b.?

OpenStudy (anonymous):

-6=3(6)-6(4)-> -6=18-24 4(6)=4+5(4)-> 24=4+20

OpenStudy (jkbo):

thank you very much

OpenStudy (anonymous):

No Problem Bro.